Hello,

Can someone help me find the volume of a sphere with a 5cm radius?

Thanks

Answer by checkley75(3666)

V=4/3PIR^3
V=4/3*22/7*5^3
V=88/21*125
V=4.19*125
V=523.75 CM^3